# Dispersion-Mediated Space-Time States

### v1

@@ V Conclusions
We have presented a general methodology for determining the spectral content of waves scattered at a uniformly moving interface between dispersive media. By combining phase-matching conditions with two conditions—group-velocity restriction and passivity—we have identified the physically allowed frequency transitions among the multiple solutions of the characteristic equations. A key feature of this methodology is that it is valid for arbitrary dispersion profiles, including regions near resonances, without relying on a Taylor expansion of the dispersion relation. It also does not require frame hopping, which makes it suitable for systems with multiple velocities or accelerated interfaces. Applying this framework to the lossless Drude model, we have showed that, in addition to the conventional nondispersive modes, a variety of dispersion-mediated space-time modes emerge. These modes have no counterpart in nondispersive scattering problems and arise from additional intersections between the dispersion curves and the phase-matching transition line. Similar behavior is observed in Lorentz media, where the higher-order characteristic equations and resonant dispersion yield an even richer set of space-time states, with up to $21$ distinct scattering regimes. These results may provide a foundational framework for analyzing realistic space-time varying systems, especially those relying on strongly dispersive materials such as epsilon-near-zero platforms in the optical regime [Boyd2020_Time_Refr_ENZ_TEM_PUB, Segev2023_Single_Cycle_TEM_PUB, Kinsey2025_ST_Knife_TEM_PUB]. By explicitly accounting for dispersion and its role in creating new frequency-transition branches, our approach offers a systematic route to dispersion-engineer space-time metamaterials.
Future works includes determining the scattering coefficients associated with each space-time state, thereby providing a complete solution to the electromagnetic scattering problem at a constant moving modulation interface in dispersive media. This task is nontrivial because it involves two distinct notions of time: an “interface time,” associated with the moving boundary, and a “memory time,” associated with the dispersive response of the medium. Dispersion is conventionally treated in the Fourier domain, whereas moving-boundary scattering problems are typically formulated in the time domain, making their consistent combination challenging. Other promising avenues for future research include extending the presented methodology to accelerated interfaces or multiple-velocity systems [Deck-Leger2019_Uni_Vel_PUB, Bahrami2025_Wedges_USTEM_PUB, Luo2025_ST_Super_Focu_USTEM_PUB].

### v2

@@ VII Conclusions
We have presented a general methodology for determining the spectral content of waves scattered at a uniformly moving interface between dispersive media. By combining phase-matching requirements with two conditions—group-velocity restriction and media passivity—we have identified the physically allowed frequency transitions among the multiple solutions of the characteristic equations. A key feature of this methodology is that it is valid for arbitrary dispersion profiles, including regions near resonances, without relying on a Taylor expansion of the dispersion relation. It also does not require frame hopping, which makes it suitable for systems with multiple velocities [Deck-Leger2019_Uni_Vel_PUB, Bahrami2025_Wedges_USTEM_PUB, Luo2025_ST_Super_Focu_USTEM_PUB] or accelerated interfaces [DeKinder2025_Scat_Chirp_ASTEM]. Applying this framework to the lossless Drude model, we have showed that, in addition to the conventional nondispersive modes, a variety of dispersion-mediated space-time modes emerge. These modes have no counterpart in nondispersive scattering problems and arise from additional intersections between the dispersion curves and the phase-matching transition line. Similar behavior is observed in Lorentz media, where the higher-order characteristic equations and resonant dispersion yield an even richer set of space-time states, with up to $21$ distinct scattering regimes. Beyond the frequency transitions, we have derived the scattering coefficients for dispersive moving interfaces in the subluminal regime and formulated a Fourier-domain approach that yields the complete electromagnetic scattering solution for both monochromatic waves and broadband pulses. This formulation consistently combines interface kinematics with the causal, frequency-dependent material response and reveals how different spectral components of a pulse may undergo distinct scattering processes, leading to strong temporal and spectral reshaping. These results may provide a foundational framework for analyzing realistic space-time varying systems, especially those relying on strongly dispersive materials such as epsilon-near-zero platforms in the optical regime [Boyd2020_Time_Refr_ENZ_TEM_PUB, Segev2023_Single_Cycle_TEM_PUB, Kinsey2025_ST_Knife_TEM_PUB]. By explicitly accounting for dispersion and its role in creating new frequency-transition branches, our approach offers a systematic route to dispersion-engineer space-time metamaterials.
While the present work provides a complete solution in regimes supporting two scattered waves, scattering scenarios involving a larger number of modes require additional boundary conditions. Identifying a systematic and physically grounded set of such conditions for general space-time varying dispersive systems remains an open problem and represents an important direction for future research. Other promising avenues for future research include extending the presented methodology to accelerated interfaces or multiple-velocity systems.
@@ Acknowledgments
Appendix A Scattering Coefficients
This section derives the electromagnetic scattering solutions at a constant-velocity moving modulation interface in the subluminal regime (region C in Fig. 1). The interface separates two isotropic, linear but dispersive media, characterized by frequency-dependent refractive indices $n_{1,2}{\left[\omega\right]}$ and impedances $\eta_{1,2}{\left[\omega\right]}$. The interface follows the trajectory $z{\left[t\right]}=v_{\text{m}}t$, parametrized with laboratory time, $t$.
We consider a one-dimensional geometry in which the interface moves along the $z$-direction, the electric field is polarized along the $x$-direction and magnetic field lies along the $y$-direction. We further consider a forward-propagating monochromatic incident wave with frequency $\omega_{\text{i}}$ in the first medium, which can be written as
$\exp\left(i\left(\beta_{1}{\left[\omega_{\text{i}}\right]}z-\omega_{\text{i}}t\right)\right)=\exp\left(i\omega_{\text{i}}\left(n_{1}{\left[\omega_{\text{i}}\right]}\frac{z}{c}-t\right)\right)\,$, (8)
where $\beta_{1}{\left[\omega_{\text{i}}\right]}=n_{1}{\left[\omega_{\text{i}}\right]}\omega_{\text{i}}/c$. Eq. (8) defines a natural traveling-wave variable $\xi=n_{1}{\left[\omega_{\text{i}}\right]}z/c-t$, with phase velocity $v_{\text{p}}=c/n_{1}{\left[\omega_{\text{i}}\right]}$. Therefore, we propose as ansatz that all scattered waves can be expressed in the general form $\psi_{i}^{\pm}{\left[\xi_{i}^{\pm}\right]}=\exp\left(i\omega_{i}^{\pm}\xi_{i}^{\pm}\right)$, where $\xi_{i}^{\pm}=\pm\left(n_{i}{\left[\omega_{i}^{\pm}\right]}z/c\mp t\right)$, $i$ denotes the medium and the superscript $\pm$ denote forward (++) propagating waves and backward (-) propagating waves.
In the subluminal regime, an incident wave generates a reflected wave ($\psi_{1}^{-}$), propagating backward in the first medium and a transmitted wave ($\psi_{2}^{+}$), propagating forward in the second medium. The electric fields may be written as
$E_{\text{i}}$ $=\psi_{\text{i}}{\left[n_{1}{\left[\omega_{\text{i}}\right]}\frac{z}{c}-t\right]}\,$, (9a)
$E_{1}^{-}$ $=\psi_{1}^{-}{\left[-\left(n_{1}{\left[\omega_{1}^{-}\right]}\frac{z}{c}+t\right)\right]}\,$, (9b)
$E_{2}^{+}$ $=\psi_{2}^{+}{\left[n_{2}{\left[\omega_{2}^{+}\right]}\frac{z}{c}-t\right]}\,$, (9c)
and the magnetic fields can be written as
$H_{\text{i}}$ $=\frac{1}{\eta_{1}{\left[\omega_{\text{i}}\right]}}\psi_{\text{i}}{\left[n_{1}{\left[\omega_{\text{i}}\right]}\frac{z}{c}-t\right]}\,$, (10a)
$H_{1}^{-}$ $=-\frac{1}{\eta_{1}{\left[\omega_{1}^{-}\right]}}\psi_{1}^{-}{\left[-\left(n_{1}{\left[\omega_{1}^{-}\right]}\frac{z}{c}+t\right)\right]}\,$, (10b)
$H_{2}^{+}$ $=\frac{1}{\eta_{2}{\left[\omega_{2}^{+}\right]}}\psi_{2}^{+}{\left[n_{2}{\left[\omega_{2}^{+}\right]}\frac{z}{c}-t\right]}\,$, (10c)
where $\psi_{\text{i}}$, $\psi_{1}^{-}$ and $\psi_{2}^{+}$ denote the incident, reflected and transmitted waves, respectively, and $\omega_{\text{i}}$, $\omega_{1}^{-}$ and $\omega_{2}^{+}$ are the corresponding frequencies  [^2] We have dropped the $(j)$ superscript used in Sec. II to distinguish different waves with same features as there is only one reflected and one transmitted wave in the subluminal regime (region C in Fig. 1).. The reflected and transmitted frequencies are determined by the general phase matching conditions [Eqs. (2)] or by their particular counterparts of Eqs. (4) for the Drude-model example [Eq. (3)].
The boundary conditions are the continuity of $E{\left[z,t\right]}-v_{\text{m}}B{\left[z,t\right]}$ and $H{\left[z,t\right]}-v_{\text{m}}D{\left[z,t\right]}$ at the moving interface. They require knowing $D{\left[z,t\right]}$ and $B{\left[z,t\right]}$, which are related to $E{\left[z,t\right]}$ and $H{\left[z,t\right]}$ via the constitutive relations. In nondispersive media, these relations are simple products: $B{\left[z,t\right]}=n\eta/cH{\left[z,t\right]}$ and $D{\left[z,t\right]}=n/\left(c\eta\right)E{\left[z,t\right]}$. In dispersive media, the constitutive relations become convolutions in time, which are cumbersome to handle  [^3] In the monochromatic case of Eq. (8), these relations become once again simple products, yet we formulate the analysis in the Fourier domain because this approach extends directly to pulses (Sec. VI), which are composed of multiple frequency components.. However, in Fourier domain, these constitutive relations reduce again to simple frequency-domain products: $\tilde{B}{\left[z,\omega\right]}=n{\left[\omega\right]}\eta{\left[\omega\right]}/c\tilde{H}{\left[z,\omega\right]}$ and $\tilde{D}{\left[z,\omega\right]}=n{\left[\omega\right]}/\left(c\eta{\left[\omega\right]}\right)\tilde{E}{\left[z,\omega\right]}$.
The time-domain boundary conditions at $z=v_{\text{m}}t$ are given by
$\left.\left(E_{\text{i}}{\left[n_{1}{\left[\omega_{\text{i}}\right]}\frac{z}{c}-t\right]}-v_{\text{m}}B_{\text{i}}{\left[n_{1}{\left[\omega_{\text{i}}\right]}\frac{z}{c}-t\right]}\right)+\left(E_{1}^{-}{\left[-\left(n_{1}{\left[\omega_{1}^{-}\right]}\frac{z}{c}+t\right)\right]}-v_{\text{m}}B_{1}^{-}{\left[-\left(n_{1}{\left[\omega_{1}^{-}\right]}\frac{z}{c}+t\right)\right]}\right)\right|_{z=v_{\text{m}}t}$
$\hskip 227.62204pt=\left.\left(E_{2}^{+}{\left[n_{2}{\left[\omega_{2}^{+}\right]}\frac{z}{c}-t\right]}-v_{\text{m}}B_{2}^{+}{\left[n_{2}{\left[\omega_{2}^{+}\right]}\frac{z}{c}-t\right]}\right)\right|_{z=v_{\text{m}}t}\,$, (11a)
$\left.\left(H_{\text{i}}{\left[n_{1}{\left[\omega_{\text{i}}\right]}\frac{z}{c}-t\right]}-v_{\text{m}}D_{\text{i}}{\left[n_{1}{\left[\omega_{\text{i}}\right]}\frac{z}{c}-t\right]}\right)+\left(H_{1}^{-}{\left[-\left(n_{1}{\left[\omega_{1}^{-}\right]}\frac{z}{c}+t\right)\right]}-v_{\text{m}}D_{1}^{-}{\left[-\left(n_{1}{\left[\omega_{1}^{-}\right]}\frac{z}{c}+t\right)\right]}\right)\right|_{z=v_{\text{m}}t}$
$\hskip 227.62204pt=\left.\left(H_{2}^{+}{\left[n_{2}{\left[\omega_{2}^{+}\right]}\frac{z}{c}-t\right]}-v_{\text{m}}D_{2}^{+}{\left[n_{2}{\left[\omega_{2}^{+}\right]}\frac{z}{c}-t\right]}\right)\right|_{z=v_{\text{m}}t}\,$, (11b)
where all fields are evaluated at the interface position. Incorporating the interface trajectory yields
$E_{\text{i}}{\left[-\left(1-n_{1}{\left[\omega_{\text{i}}\right]}\frac{v_{\text{m}}}{c}\right)t\right]}-v_{\text{m}}B_{\text{i}}{\left[-\left(1-n_{1}{\left[\omega_{\text{i}}\right]}\frac{v_{\text{m}}}{c}\right)t\right]}+E_{1}^{-}{\left[-\left(1+n_{1}{\left[\omega_{1}^{-}\right]}\frac{v_{\text{m}}}{c}\right)t\right]}-v_{\text{m}}B_{1}^{-}{\left[-\left(1+n_{1}{\left[\omega_{1}^{-}\right]}\frac{v_{\text{m}}}{c}\right)t\right]}$
$\hskip 199.16928pt=E_{2}^{+}{\left[-\left(1-n_{2}{\left[\omega_{2}^{+}\right]}\frac{v_{\text{m}}}{c}\right)t\right]}-v_{\text{m}}B_{2}^{+}{\left[-\left(1-n_{2}{\left[\omega_{2}^{+}\right]}\frac{v_{\text{m}}}{c}\right)t\right]}\,$, (12a)
$H_{\text{i}}{\left[-\left(1-n_{1}{\left[\omega_{\text{i}}\right]}\frac{v_{\text{m}}}{c}\right)t\right]}-v_{\text{m}}D_{\text{i}}{\left[-\left(1-n_{1}{\left[\omega_{\text{i}}\right]}\frac{v_{\text{m}}}{c}\right)t\right]}+H_{1}^{-}{\left[-\left(1+n_{1}{\left[\omega_{1}^{-}\right]}\frac{v_{\text{m}}}{c}\right)t\right]}-v_{\text{m}}D_{1}^{-}{\left[-\left(1+n_{1}{\left[\omega_{1}^{-}\right]}\frac{v_{\text{m}}}{c}\right)t\right]}$
$\hskip 199.16928pt=H_{2}^{+}{\left[-\left(1-n_{2}{\left[\omega_{2}^{+}\right]}\frac{v_{\text{m}}}{c}\right)t\right]}-v_{\text{m}}D_{2}^{+}{\left[-\left(1-n_{2}{\left[\omega_{2}^{+}\right]}\frac{v_{\text{m}}}{c}\right)t\right]}\,.$ (12b)
Now comes the central idea of the proposed method. We Fourier-transform Eqs. (12), while including the frequency-domain constitutive relations, inserting the general waveforms Eqs. (9) and Eqs. (10), and using the Fourier scaling property
$\mathcal{F}\left\{f{\left[at\right]}\right\}{\left[\omega\right]}=\frac{1}{\left|a\right|}\tilde{f}{\left[\frac{\omega}{a}\right]}\,.$ (13)
This leads to the following simple system of equations for the reflected and transmitted waves:
$\tilde{\psi}_{\text{i}}\left[-\frac{\omega}{1-n_{1}{\left[\omega_{\text{i}}\right]}v_{\text{m}}/c}\right]+\tilde{\psi}_{1}^{-}\left[-\frac{\omega}{1+n_{1}{\left[\omega_{1}^{-}\right]}v_{\text{m}}/c}\right]$ $=\tilde{\psi}_{2}^{+}\left[-\frac{\omega}{1-n_{2}{\left[\omega_{2}^{+}\right]}v_{\text{m}}/c}\right]\,$, (14a)
$\frac{1}{\eta_{1}{\left[\omega_{\text{i}}\right]}}\tilde{\psi}_{\text{i}}\left[-\frac{\omega}{1-n_{1}{\left[\omega_{\text{i}}\right]}v_{\text{m}}/c}\right]-\frac{1}{\eta_{1}{\left[\omega_{1}^{-}\right]}}\tilde{\psi}_{1}^{-}\left[-\frac{\omega}{1+n_{1}{\left[\omega_{1}^{-}\right]}v_{\text{m}}/c}\right]$ $=\frac{1}{\eta_{2}{\left[\omega_{2}^{+}\right]}}\tilde{\psi}_{2}^{+}\left[-\frac{\omega}{1-n_{2}{\left[\omega_{2}^{+}\right]}v_{\text{m}}/c}\right]\,.$ (14b)
Solving Eqs. (14) for the reflected and transmitted waves yields
$\tilde{\psi}_{1}^{-}\left[-\frac{\omega}{1+n_{1}{\left[\omega_{1}^{-}\right]}v_{\text{m}}/c}\right]$ $=\frac{\eta_{1}{\left[\omega_{1}^{-}\right]}}{\eta_{1}{\left[\omega_{\text{i}}\right]}}\frac{\eta_{2}{\left[\omega_{2}^{+}\right]}-\eta_{1}{\left[\omega_{\text{i}}\right]}}{\eta_{2}{\left[\omega_{2}^{+}\right]}+\eta_{1}{\left[\omega_{1}^{-}\right]}}\tilde{\psi}_{\text{i}}\left[-\frac{\omega}{1-n_{1}{\left[\omega_{\text{i}}\right]}v_{\text{m}}/c}\right]\,$, (15a)
$\tilde{\psi}_{2}^{+}\left[-\frac{\omega}{1-n_{2}{\left[\omega_{2}^{+}\right]}v_{\text{m}}/c}\right]$ $=\frac{\eta_{2}{\left[\omega_{2}^{+}\right]}}{\eta_{1}{\left[\omega_{\text{i}}\right]}}\frac{\eta_{1}{\left[\omega_{1}^{-}\right]}+\eta_{1}{\left[\omega_{\text{i}}\right]}}{\eta_{1}{\left[\omega_{1}^{-}\right]}+\eta_{2}{\left[\omega_{2}^{+}\right]}}\tilde{\psi}_{\text{i}}\left[-\frac{\omega}{1-n_{1}{\left[\omega_{\text{i}}\right]}v_{\text{m}}/c}\right]\,.$ (15b)
To express the scattered fields again as functions of the usual frequency variable, $\omega$, we finally introduce spectral variables that absorb the Doppler prefactors appearing in Eqs. (15),
$\omega$ $\mapsto-\left(1+n_{1}{\left[\omega_{1}^{-}\right]}\frac{v_{\text{m}}}{c}\right)\omega\,$, (16a)
$\omega$ $\mapsto-\left(1-n_{2}{\left[\omega_{2}^{+}\right]}\frac{v_{\text{m}}}{c}\right)\omega\,.$ (16b)
Inserting Eqs. (16) into Eqs. (15) yields the final solutions
$\tilde{\psi}_{1}^{-}\left[\omega\right]$ $=\frac{\eta_{1}{\left[\omega_{1}^{-}\right]}}{\eta_{1}{\left[\omega_{\text{i}}\right]}}\frac{\eta_{2}{\left[\omega_{2}^{+}\right]}-\eta_{1}{\left[\omega_{\text{i}}\right]}}{\eta_{2}{\left[\omega_{2}^{+}\right]}+\eta_{1}{\left[\omega_{1}^{-}\right]}}\tilde{\psi}_{\text{i}}\left[\frac{1+n_{1}{\left[\omega_{1}^{-}\right]}v_{\text{m}}/c}{1-n_{1}{\left[\omega_{\text{i}}\right]}v_{\text{m}}/c}\omega\right]\,$, (17a)
$\tilde{\psi}_{2}^{+}\left[\omega\right]$ $=\frac{\eta_{2}{\left[\omega_{2}^{+}\right]}}{\eta_{1}{\left[\omega_{\text{i}}\right]}}\frac{\eta_{1}{\left[\omega_{1}^{-}\right]}+\eta_{1}{\left[\omega_{\text{i}}\right]}}{\eta_{2}{\left[\omega_{2}^{+}\right]}+\eta_{1}{\left[\omega_{1}^{-}\right]}}\tilde{\psi}_{\text{i}}\left[\frac{1-n_{2}{\left[\omega_{2}^{+}\right]}v_{\text{m}}/c}{1-n_{1}{\left[\omega_{\text{i}}\right]}v_{\text{m}}/c}\omega\right]\,.$ (17b)